# On the Impacts of Contexts on Repository-Level Code Generation

### v3

@@ 4.3.2 Coverage Enhancement
Weak unit tests may allow incorrect implementations to pass Liu et al. (2023a). To address this, we propose a strategy for enhancing test coverage using LLMs. Given the complexity of this task, requiring a strong understanding of code, we utilize GPT-3.5 to improve the quality of test cases. We provide GPT-3.5 with three prompts (see Appendix C) to handle challenging scenarios, including edge and corner cases. The initial generated tests serve as few-shot examples. We then extract the newly generated tests and ensure their correctness using our methodology from Section 4.3.1. Table 1 shows a line coverage improvement of about 4% after enhancement, reaching 96.25%. The performance gap has also increased to over 5% (Appendix B), indicating greater robustness in the evaluation.
@@ 5.1 Data Formatting
Figure 2 illustrates the input data format used in RepoExec. We retain import information and append dependencies in the order presented in the import statements. The target function signature and natural language description are positioned at the end of the input prompt. We propose three prompt types with varying context lengths to test the reasoning capability of CodeLLMs in leveraging contexts for repo-level code generation:
@@
We follow Muennighoff et al. 2023 to define 2 types of prompt formats in our evaluation of RepoExec across LLMs: (1) BasePrompt, which concatenates all contexts with the target functions (Figure 2), and (2) InstructPrompt, which includes specific instructions for the LLMs to follow, utilizing two variations as input formats (further details and examples in Appendix D).
@@ 5.2 Dataset Stastistic
Dataset #Samples Context TC LC (%) CoNaLA Yin et al. (2018) 500 - CONCODE Iyer et al. (2018) 2,000 - HumanEval Chen et al. (2021) 164 ✓, H 99.43 MBPP Austin et al. (2021) 974 ✓, H 98.48 RepoCoder Zhang et al. (2023b) 373 ✓, P - CrossCodeEval Ding et al. (2023) 2,665 - CoCoMIC Ding et al. (2024) 6,888 - DevEval Li et al. (2024) 1,874 ✓, P - RepoExec 355 ✓, A 92.46    + coverage-enhancement 96.25
Table 1: The comparison between popular code generation benchmarks and RepoExec. For test case, we denote H: Human annotated, P: Pre-existing, A: Automated. TC: Test Cases. LC: Line Coverage
Table 1 compares the details of RepoExec with those of existing code generation benchmarks. Benchmarks that exclude execution-based evaluation Yin et al. (2018); Iyer et al. (2018); Ding et al. (2023, 2024) may gather substantial amounts of data; however, they are inadequate for assessing the quality of the generated code. For HumanEval and MBPP, the majority of problems involve standalone functions, which do not reflect real-world scenarios. Besides, benchmarks that rely on human-annotated and pre-existing test cases Chen et al. (2021); Austin et al. (2021); Zhang et al. (2023a); Li et al. (2024) are challenging to scale and control the test coverage. Finally, repository-context benchmarks (RepoCoder Zhang et al. (2023a), CrossCodeEval Ding et al. (2023), CoCoMIC Ding et al. (2024), and DevEval Li et al. (2024)) primarily focus on evaluating retrieval modules. For example, RepoCoder introduces a retriever using a sparse bag-of-words model and provides the effectiveness of this retriever. Similarly, CrossCodeEval experiments and reports performance using various retrievers such as BM25, UniXCoder, and OpenAI ada. CoCoMic proposed CCFINDER to retrieve cross-file context. DevEval is the closest to our work; however, they compare the performance of the generation model based on different given types of context, which can also align with the different contexts provided by different retrievers. In contrast, our approach emphasizes the generation module’s ability to understand and utilize human-provided dependency contexts.
Dataset attributes: Table 2 outlines the characteristics of RepoExec, including the total number of examples, the average number of test cases per problem, and the number of tokens in prompts and solutions.
Table 2: Dataset attributes with different levels of contexts (Full, Medium, Small). Cross-file refers to the percentage of problems that involve cross-file dependencies. Tree-sitter is used for tokenization. AVG LC: Average Line Coverage.
@@ 6.1 Evaluation Results
Model Full context Medium context Small context pass@1 pass@5 DIR pass@1 pass@5 DIR pass@1 pass@5 DIR BasePrompt Pre-models CodeLlama-34b-Python Roziere et al. (2023) 42.93 49.54 68.85 35.92 42.95 58.15 39.80 45.79 64.23 CodeLlama-13b-Python Roziere et al. (2023) 38.65 43.24 62.26 32.96 38.33 56.38 35.66 42.41 62.67 StarCoder Li et al. (2023) 28.08 33.95 58.75 22.54 31.83 50.74 25.54 31.45 56.67 StarCoder2-15b Lozhkov et al. (2024) 27.77 32.60 60.57 18.70 23.28 39.28 23.27 29.78 53.49 Mixtral-8x7B-v0.1 Jiang et al. (2024) 22.82 29.14 55.90 19.38 25.25 47.22 20.54 26.30 53.40 Phi-2 Javaheripi et al. (2023) 19.04 24.56 48.22 14.54 20.34 40.85 14.82 20.69 44.54 Phi-1 Gunasekar et al. (2023) 14.99 18.38 43.17 12.48 15.42 37.45 12.54 15.96 38.75 Inst-models WizardCoder-Python-13B-V1.0 Luo et al. (2023) 34.31 40.06 62.90 30.99 36.75 59.50 32.54 38.34 64.67 Phind-CodeLlama-34B-v2 30.08 33.49 59.47 25.25 29.40 50.73 27.55 31.85 58.93 CodeLlama-13b-Instruct Roziere et al. (2023) 28.56 32.67 57.09 26.25 30.72 49.48 26.73 33.50 54.53 GPT-3.5 27.27 37.69 63.59 23.15 33.94 52.79 22.59 33.63 55.22 DeepSeek-Coder-7b-Instruct Guo et al. (2024) 25.18 29.91 58.50 20.23 26.02 45.76 22.20 27.74 56.69 Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024) 23.41 28.71 59.83 19.04 24.83 52.75 20.45 25.84 58.17 InstructPrompt Inst-models WizardCoder-Python-13B-V1.0 26.20 30.68 67.32 24.56 30.25 67.54 24.70 29.56 68.34 CodeLlama-13b-Instruct 25.66 30.82 62.04 27.44 34.11 63.33 26.73 32.43 64.85 GPT-3.5 23.82 39.10 40.55 19.62 36.03 37.48 19.00 34.00 35.28 Mixtral-8x7B-Instruct-v0.1 18.11 23.04 67.73 18.54 23.12 69.66 15.38 20.61 68.86
Table 3: Pass@k (k$=1$ and $5$) and DIR results of various LLMs on RepoExec. Bold scores indicate the highest values, while Underlined scores represent the second highest. Pre- and Inst- denote Pretrained and Instruction-tuned, respectively.
We evaluated 13 CodeLLMs on RepoExec and presented the results (pass@1, pass@5, and DIR) in Table 3. We use nucleus sampling with temperature set to 0.2, top-p to 0.95, and 10 outputs generated for all models. The results show that retaining the full context of dependencies yields the best performance across all models. Surprisingly, Small-size context proves to be more effective than Medium-size context, which we attribute to the context’s input format using BasePrompt, potentially misleading the model into interpreting the dependency functions as few-shot examples. Using the Small-size context results in a fair decrease in performance compared to the Full context while effectively reducing the input length by half.
Codellama-34b-Python achieves the highest pass@1 rate among the models considered, and foundation models demonstrate greater effectiveness compared to instruction-tuned models on our dataset. However, our analysis reveals inherent limitations in both types of models:
@@
    Pretrained LLMs produce code that functions correctly but do not effectively reuse the provided dependencies, occasionally duplicating the implementation from the given context. This leads to redundancy and potentially creates technical debt or code smell issues (Maldonado and Shihab, 2015; Sierra et al., 2019; Rasool and Arshad, 2015; Santos et al., 2018).
These issues can lead to a low-quality codebase, requiring substantial human effort for reviewing and fixing, which may even exceed the effort needed to write the code from scratch. Details and examples are discussed further in Appendix E.
6.2 Generation with Multi-round Debugging
In this section, we examine the models’ self-refinement capabilities in enhancing generation performance. We provide the models with error output logs and ask them to fix the errors. We experiment with WizardCoder, GPT3.5, and CodeLlama-13b-Python. The number of rounds to debug is set to 3 and the input template is presented in Appendix F. In this experiment, we employ a greedy search algorithm to generate only a single output.
@@ 6.3 Instruction-tuning with Code Dependencies
While the multi-round debugging experiment has demonstrated effectiveness in leveraging given dependencies to provide correct solutions (Section 6.2), it requires a strong model to generate good test cases and can be time-consuming due to the repeated generation and execution of code and test cases. To address these challenges, we propose an instruction-tuning dataset for fine-tuning base LLMs. We collected the 1,555 most-starred repositories from 2018 onward and extracted functions with their corresponding dependencies, following the procedure outlined in Section 4.2. We obtained 154,818 functions, of which 57,746 samples include docstrings. From this set, we utilized 50K samples with docstrings and applied instruction prompts, while 80K samples adhered to the raw code format (Full context). Recognizing the potential of the Small context format, we allocated the remaining 20K samples to follow this structure. We fine-tuned Phi-2, StarCoder, StarCoder2, and CodeLlama-13b-Python models for 5 epochs with LoRa Hu et al. (2021) and used 10% of the training data as the validation set to select the best checkpoint.
Table 5 illustrates the efficacy of our training data. All 4 models demonstrate improvements in both Pass@1 and DIR after instruction tuning. Specifically, there is a slight increase of around 1% in Pass@k for all models, while DIR shows a significant improvement, reaching the highest scores (over 70%) compared to other models after tuning with our dataset. Notably, the performance improves significantly with the use of small context, achieving results comparable to those obtained with full context. This offers the potential for more efficient processing and reducing computational costs, allowing additional space to integrate various other types of context. In summary, our proposed instruction-tuning method can improve the model’s ability to both reuse dependencies and ensure the functional correctness of the output. Although self-refinement through multi-round debugging (Section 6.2) demonstrates greater effectiveness, instruction-tuning models only utilize single-turn generation, making them more efficient in practice. To facilitate open research in this domain, we will publicly release this dataset.
@@ 7 Conclusion
We introduce an evaluation approach for repository-level code generation that rethinks the limitations of previous methods by assessing not only functional correctness but also dependency utilization to ensure the quality of the generated code. By presenting RepoExec, a novel Python code generation benchmark at the repository level with executable capabilities, designed to evaluate the alignment of generated code with developer intent and its correctness. Our experiments reveal that while pretrained language models (LLMs) excel in functional correctness, instruction-tuned models demonstrate proficiency in utilizing dependencies and debugging. However, existing models struggle to effectively reuse provided dependencies, potentially leading to technical debt and code smell issues. We provide a comprehensive evaluation of how different LLMs leverage code dependencies when generating code, and our findings show that the size of the context significantly impacts the final results.
We also introduce an instruction-tuning dataset that enhances dependency invocation accuracy and output correctness, even with limited context. This approach allows for the integration of additional context types and mitigates large token length constraints. Our contributions establish a foundation for future research in code generation, providing valuable evaluation techniques to drive the development of more capable and reliable models.
@@ 8 Limitations
RepoExec is constructed for evaluating models based on existing works. However, this approach can potentially lead to data leakage, especially when modern models are trained on similar datasets. If the benchmark relies heavily on known works, there’s a risk that the model may inadvertently learn from specific patterns or features present in those works, compromising its generalization ability. Despite this concern, experimental results demonstrate that even modern models struggle to handle the challenges posed by RepoExec, indicating that the benchmark remains a valuable tool for assessing model performance. Besides, several models, including StarCoder2 and DeepSeek-Coder, have been pretrained using repository-level context. However, these models typically concatenate the contents of multiple files in a repository without filtering out irrelevant information or considering the selection of dependencies. This helps our dataset distinguish from the pretraining datasets of these models, thereby helping to mitigate the data leakage issue. However, employing new repositories can help mitigate these phenomena.
In this work, we only consider one level of dependency context, specifically the dependencies directly called from the target function. While this simplification facilitates manageable analysis and model development, it may not fully capture the valuable context necessary for leveraging models effectively. However, incorporating deep dependencies could significantly extend the input length, posing challenges in managing long context inputs and potentially exceeding the maximum input length. Our approach has demonstrated promising outcomes with the small-size context version, creating opportunities for integrating additional input context. Future research could explore incorporating multiple levels of dependencies, creating a more comprehensive graph that includes transitive dependencies, indirect calls, and broader contextual information. By doing so, we could enhance the model’s understanding of code interactions and improve its ability to handle intricate software execution scenarios.
@@
Appendix A Evaluation Metrics: Match-Based vs. Execution-Based

### v4

@@ 4.2.2 Coverage Enhancement
Weak unit tests may allow incorrect implementations to pass Liu et al. (2023a). To address this, we propose a strategy for enhancing test coverage using LLMs. Given the complexity of this task, requiring a strong understanding of code, we utilize GPT-3.5 to improve the quality of test cases. We provide GPT-3.5 with three prompts (see Appendix D) to handle challenging scenarios, including edge and corner cases. The initial generated tests serve as few-shot examples. We then extract the newly generated tests and ensure their correctness using our methodology from Section 4.2.1. Table 3 shows a line coverage improvement of about 4% after enhancement, reaching 96.25%. The performance gap has also increased to over 5% (Appendix C), indicating greater robustness.
@@ 5 Data Characteristics
Pipeline RepoContext Test Test Mining Generation Enhancement RepoCoder R2E CodeBenchGen RepoExec
Table 2: Comparison of data construction pipelines across different repo-level code generation benchmarks.
@@ 5.1 Data Formatting
Figure 2 illustrates the input data format used in RepoExec. We retain import information and append dependencies in the order presented in the import statements, placing the target function signature and description at the end of the prompt. To evaluate the reasoning capability of CodeLLMs in repo-level code generation, we propose three prompt types with varying context lengths:
@@
We follow Muennighoff et al. 2023 to define 2 types of prompt formats in our evaluation of RepoExec across LLMs: (1) BasePrompt, which concatenates all contexts with the target functions (Figure 2), and (2) InstructPrompt, which includes specific instructions for the LLMs to follow, utilizing two variations as input formats (further details and examples in Appendix E).
@@ 5.2 Dataset Stastistic
Dataset #Samples RC TC LC (%) CoNaLA Yin et al. (2018) 500 - CONCODE Iyer et al. (2018) 2,000 - HumanEval Chen et al. (2021) 164 ✓, H 99.43 MBPP Austin et al. (2021) 974 ✓, H 98.48 RepoCoder Zhang et al. (2023b) 373 ✓, P - CrossCodeEval Ding et al. (2023) 2,665 - CoCoMIC Ding et al. (2024) 6,888 - DevEval Li et al. (2024) 1,874 ✓, P - RepoExec 355 ✓, A 92.46    + coverage-enhancement 96.25
Table 3: The comparison between popular code generation benchmarks and RepoExec. For test case, we denote H: Human annotated, P: Pre-existing, A: Automated. RC: Repo-context utilization. TC: Test Cases. LC: Line Coverage
Model Full context Medium context Small context pass@1 pass@5 DIR pass@1 pass@5 DIR pass@1 pass@5 DIR BasePrompt Pre-models CodeLlama-13b-Python Roziere et al. (2023) 38.65 43.24 62.26 32.96 38.33 56.38 35.66 42.41 62.67 StarCoder Li et al. (2023) 28.08 33.95 58.75 22.54 31.83 50.74 25.54 31.45 56.67 StarCoder2-15b Lozhkov et al. (2024) 27.77 32.60 60.57 18.70 23.28 39.28 23.27 29.78 53.49 Mixtral-8x7B-v0.1 Jiang et al. (2024) 22.82 29.14 55.90 19.38 25.25 47.22 20.54 26.30 53.40 Phi-2 Javaheripi et al. (2023) 19.04 24.56 48.22 14.54 20.34 40.85 14.82 20.69 44.54 Phi-1 Gunasekar et al. (2023) 14.99 18.38 43.17 12.48 15.42 37.45 12.54 15.96 38.75 Inst-models DeepSeek-R1 Guo et al. (2025) 42.57 - 70.86 - - - - - - DeepSeek-V3 Liu et al. (2024a) 42.00 - 80.35 - - - - - - Llama 3.1-405B-Instruct Dubey et al. (2024) 34.86 - 78.81 - - - - - - GPT-4o 37.14 - 81.43 - - - - - - GPT-4o-mini 30.29 - 74.75 - - - - - - CodeLlama-34b-Python Roziere et al. (2023) 40.93 49.54 68.85 35.92 42.95 58.15 39.80 45.79 64.23 WizardCoder-Python-13B-V1.0 Luo et al. (2023) 34.31 40.06 62.90 30.99 36.75 59.50 32.54 38.34 64.67 Phind-CodeLlama-34B-v2 30.08 33.49 59.47 25.25 29.40 50.73 27.55 31.85 58.93 CodeLlama-13b-Instruct Roziere et al. (2023) 28.56 32.67 57.09 26.25 30.72 49.48 26.73 33.50 54.53 GPT-3.5 27.27 37.69 63.59 23.15 33.94 52.79 22.59 33.63 55.22 DeepSeek-Coder-7b-Instruct Guo et al. (2024) 25.18 29.91 58.50 20.23 26.02 45.76 22.20 27.74 56.69 Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024) 23.41 28.71 59.83 19.04 24.83 52.75 20.45 25.84 58.17 InstructPrompt Inst-models DeepSeek-R1 Guo et al. (2025) 39.71 - 53.78 - - - - - - DeepSeek-V3 Liu et al. (2024a) 41.71 - 62.58 - - - - - - Llama 3.1-405B-Instruct Dubey et al. (2024) 39.43 - 73.26 - - - - - - GPT-4o 38.00 - 68.15 - - - - - - GPT-4o-mini 32.29 - 53.34 - - - - - - WizardCoder-Python-13B-V1.0 26.20 30.68 67.32 24.56 30.25 67.54 24.70 29.56 68.34 CodeLlama-13b-Instruct 25.66 30.82 62.04 27.44 34.11 63.33 26.73 32.43 64.85 GPT-3.5 23.82 39.10 40.55 19.62 36.03 37.48 19.00 34.00 35.28 Mixtral-8x7B-Instruct-v0.1 18.11 23.04 67.73 18.54 23.12 69.66 15.38 20.61 68.86
Table 4: Pass@k and DIR results of various LLMs on RepoExec. Bold scores indicate the highest, while Underlined scores denote the second highest. Pre- and Inst- denote Pretrained and Instruction-tuned, respectively. More results are presented in Table 8.
@@
Table 3 compares the details of RepoExec with existing code generation benchmarks. Benchmarks that exclude execution-based evaluation Yin et al. (2018); Iyer et al. (2018); Ding et al. (2023, 2024) may gather substantial amounts of data; however, they are inadequate for assessing the quality of the generated code. For HumanEval and MBPP, the majority of problems involve standalone functions, which do not reflect real-world scenarios. Besides, benchmarks that rely on human-annotated and pre-existing test cases Chen et al. (2021); Austin et al. (2021); Zhang et al. (2023a); Li et al. (2024) are challenging to scale and control the test coverage. Finally, repository-context benchmarks Zhang et al. (2023a); Ding et al. (2023, 2024); Li et al. (2024) primarily focus on investigating retrieval modules. For example, RepoCoder analyzes a retriever using a sparse bag-of-words model. Similarly, CrossCodeEval experiments and reports performance using various retrievers such as BM25, UniXCoder, and OpenAI ada. CoCoMic proposed CCFINDER to retrieve cross-file context. DevEval is the closest to our work; however, they compare the performance of the generation model based on different given types of file-level context, which can also align with the different contexts provided by different retrievers. In contrast, our approach emphasizes the generation module’s ability to understand and utilize human-provided dependency contexts.
Comparison to Existing Data Pipeline:
Several studies, including RepoCoder (Zhang et al., 2023a), R2E Jain et al. (2024), and CodeBenchGen Xie et al. (2024), have introduced data pipelines for repository-level code generation. Table 2 compares our pipeline with these approaches across three key aspects: RepoContext mining, Test Generation, and Test Enhancement. These components are crucial for robust evaluation, as missing any can compromise quality. RepoCoder extracts existing test cases but lacks enhancement, limiting scalability. CodeBenchGen uses LLMs to synthesize function contexts, reducing practicality. Lastly, R2E is the most closely related to our approach, however, it similarly neglects test case enhancement like RepoCoder, which can undermine evaluation robustness (as detailed in Appendix C).
Dataset attributes:
Table 1 outlines the characteristics of RepoExec, including the total number of examples, the average number of test cases, and the number of tokens in prompts and solutions.
@@ 6.1 Evaluation Results
We evaluated 18 CodeLLMs on RepoExec and presented the results (pass@1, pass@5, and DIR) in Table 4. We use nucleus sampling with temperature set to 0.2, top-p to 0.95, and 10 outputs generated for all models. For evaluating super-large models requiring paid API access, we utilize greedy decoding and report pass@1 and DIR metrics under the Full Context setting. The results show that retaining the full context of dependencies yields the best performance across all models. Surprisingly, Small-size context proves to be more effective than Medium-size context, which we attribute to the context’s input format using BasePrompt, potentially misleading the model into interpreting the dependency functions as few-shot examples. Using the Small-size context results in a fair decrease in performance compared to the Full context while effectively reducing the input length.
DeepSeek-R3 achieves the highest pass@1 rate among the evaluated models. Additionally, BasePrompt and pretrained models exhibit superior effectiveness over instruction-tuned models in terms of functional correctness (pass@k) when comparing models of similar size on RepoExec. However, our analysis reveals limitations in both types of models. Additional discussions and examples are provided in Appendix F.
@@
    Pretrained LLMs generate functionally correct code but often fail to effectively utilize the provided dependencies, occasionally reimplementing dependencies already present in the given context. This leads to redundancy and potentially creates technical debt or code smell issues (Maldonado and Shihab, 2015; Sierra et al., 2019; Santos et al., 2018; Hai et al., 2024a).
These issues can lead to a low-quality codebase, requiring substantial human effort for reviewing and fixing, which may even exceed the effort needed to write the code from scratch. Moreover, the performance of current advanced LLMs highlights that DeepSeek models, particularly DeepSeek-R1 and DeepSeek-V3, demonstrate strong overall capabilities, effectively balancing accuracy and dependency management. Among smaller models, GPT-4o-mini achieves the highest DIR (74.75) but underperforms in pass@1 (30.29), emphasizing it often invokes dependencies incorrectly despite its emphasis on handling them. Models with high pass@k and DIR scores, such as DeepSeek-V3 (80.35) and Llama 3.1 (73.26), excel in accurately invoking dependencies, which is crucial for real-world code generation tasks. Finally, RepoExec remains challenging to existing models, as evidenced by consistently low performance, underscoring the need for more advanced systems.
@@
6.2 Generation with Multi-round Debugging
In this section, we examine the models’ self-refinement capabilities in enhancing generation performance. We provide the models with error output logs and ask them to fix the errors. We experiment with WizardCoder, GPT3.5, and CodeLlama-13b-Python. The number of rounds to debug is set to 3 and the input template is presented in Appendix G. In this experiment, we employ a greedy search algorithm to generate only a single output.
@@ 6.3 Instruction-tuning with Code Dependencies
While the multi-round debugging experiment has demonstrated effectiveness in leveraging given dependencies to provide correct solutions (Section 6.2), it requires a strong model to generate good test cases and can be time-consuming due to the repeated generation and execution of code and test cases. To address these challenges, we propose an instruction-tuning dataset for fine-tuning base LLMs. We collected the 1,555 most-starred repositories from 2018 onward and extracted functions with their corresponding dependencies, following the procedure outlined in Section 4.1. We obtained 154,818 functions, of which 57,746 samples include docstrings. We use 50K samples with docstrings and applied instruction prompts, while 80K samples adhered to the raw code format (Full context). Recognizing the potential of the Small context format, we allocated the remaining 20K samples to follow this structure. We fine-tuned Phi-2, StarCoder, StarCoder2, and CodeLlama-13b-Python models for 5 epochs with LoRa Hu et al. (2021) and used 10% of the training data as the validation set to select the best checkpoint.
Table 6 illustrates the efficacy of our training data. All 4 models demonstrate improvements in both Pass@1 and DIR after instruction tuning. Specifically, there is a slight increase of around 1% in Pass@k for all models, while DIR shows a significant improvement, reaching the highest scores (over 70%) compared to other models after tuning with our dataset. Notably, performance improves significantly with small context, matching results from full context, enabling more efficient processing and reducing computational costs. In summary, our instruction-tuning method enhances the model’s ability to utilize dependencies and ensure functional correctness. While multi-round debugging (Section 6.2) is more effective, instruction-tuned models rely on single-turn generation, making them more practical and efficient. To facilitate open research in this domain, we will publicly release this dataset.
@@ 7 Conclusion
We propose an evaluation approach for repository-level code generation that rethinks the limitations of prior methods by assessing not only the functional correctness aspect but also dependency utilization to ensure code quality. We introduce RepoExec, a novel Python code generation benchmark with executable capabilities, designed to evaluate the alignment of generated code with developer intent and correctness. Our experiments show that while pretrained LLMs excel in functional correctness, instruction-tuned models perform better in utilizing dependencies and debugging. However, existing models struggle to reuse provided dependencies, risking technical debt and code smells. We provide a comprehensive analysis of how LLMs leverage dependencies, revealing that context information significantly influences the results. Besides, we also introduce an instruction-tuning dataset that enhances dependency invocation accuracy and output correctness, even with limited context. Our contributions establish a foundation for future research in code generation, providing valuable evaluation techniques to drive the development of more capable and reliable models.
@@ 8 Limitations
In this work, we currently consider one level of dependency context, specifically the dependencies directly called from the target function. While this simplification facilitates manageable analysis and model development, it may not fully capture the valuable context necessary for leveraging models effectively. However, incorporating deep dependencies could significantly extend the input length, posing challenges in managing long context inputs and potentially exceeding the maximum input length. Our approach has demonstrated promising outcomes with the small-size context version, creating opportunities for integrating additional input context. Future research could explore incorporating multiple levels of dependencies, creating a more comprehensive graph that includes transitive dependencies, indirect calls, and broader contextual information. By doing so, we could enhance the model’s understanding of code interactions and improve its ability to handle intricate software execution scenarios.
@@
Appendix A Data Source
Creating an executable benchmark within repository contexts poses significant challenges due to intricate setup requirements and the often inadequate installation instructions provided in repositories. Therefore, we select repositories that can be automatically set up using pipreqs to auto-detect the required packages. Additionally, to effectively utilize the test generation module of our pipeline, we select repositories that align with the literature on unit test generation. As a result, RepoExec, is based on repositories from the unit test generation studies (Schäfer et al., 2023; Lukasczyk et al., 2023; Lemieux et al., 2023; Liu et al., 2024c), which provide executable environments and robust test generation capabilities.
Some might claim that our dataset is constructed for based on existing works, thus, this approach can potentially lead to data leakage, especially when modern models are trained on similar datasets. If the benchmark relies heavily on known works, there’s a risk that the model may inadvertently learn from specific patterns or features present in those works, compromising its generalization ability. Despite this concern, experimental results demonstrate that even modern models struggle to handle the challenges posed by RepoExec, indicating that the benchmark remains a valuable tool for assessing model performance. Besides, several models, including StarCoder2 and DeepSeek-Coder, have been pretrained using repository-level context. However, these models typically concatenate the contents of multiple files in a repository without filtering out irrelevant information or considering the selection of dependencies. This helps our dataset distinguish from the pretraining datasets of these models, thereby helping to mitigate the data leakage issue. To further measure the data leakage level compare to other benchmarks, we follow BenBench Xu et al. (2024) to report the perplexity and n-gram metrics, against three datasets—HumanEval, RepoCoder, and DevEval—analyzing their levels of data leakage when tested on three powerful LLMs. As indicated in Table 7, while HumanEval exhibits a significant level of data leakage, as expected, RepoExec demonstrates notably low levels of leakage. This is reflected in its higher perplexity and reduced n-gram overlap in most cases. The extent of data leakage in RepoExec is comparable to that of RepoCoder, a recently developed benchmark based on repositories created after January 1, 2022, to mitigate the data leakage problem. Consequently, the leakage issue in RepoExec is minimal, aligning with benchmarks like RepoCoder that utilize the latest repositories. We believe that repository-level code generation could be a valuable contribution to more robust and high-quality evaluation for advanced agent systems and dynamic environments (Huang et al., 2023; Phan et al., 2024b; Yadav et al., 2023; Hai et al., 2024b; Tran et al., 2024). By assessing models in a more comprehensive coding context, this approach enables a deeper evaluation of their ability to understand, integrate, and adapt code components effectively.
Model perplexity n-gram HumanEval RepoCoder DevEval RepoExec HumanEval RepoCoder DevEval RepoExec Mistral-7B-v0.3 2.10 177.99 170.67 120.77 0.35 0.30 0.26 0.18 CodeLlama-7B-Python 1.96 90.71 32.14 72.77 0.43 0.30 0.43 0.30 CodeLlama-13B-Python 1.96 72.66 23.83 126.51 0.48 0.31 0.41 0.31
Table 7: Comparison of data leakage level between RepoExec and other benchmarks.
Appendix B Evaluation Metrics: Match-Based vs. Execution-Based